# Defeaturing.bib

### v1

@@ 3 Error estimator via equilibrated flux reconstruction
We now analyze the properties of $\bm{\sigma}_{h}$ in different scenarios, in particular when $\mathcal{F}^{\star}$ is or not an empty set. For any $\bm{a}\in\mathscr{N}_{h}$ we denote by $\overline{q_{h}}^{\omega_{\bm{a}}}$ the average of $q_{h}\in Q_{h}({\omega_{\bm{a}}})$ on ${\omega_{\bm{a}}}$. Setting $c_{\bm{a}}=-\overline{q_{h}}^{\omega_{\bm{a}}}$, we have that $q_{h}+c_{\bm{a}}\in Q_{h}^{\bm{a}}$. The second equation in (23) can hence be rewritten as
$b_{\bm{a}}^{p}(\bm{\bm{\sigma}_{h}^{\bm{a}}},q_{h}+c_{\bm{a}})=R_{\bm{a}}^{p}(q_{h}+c_{\bm{a}})\qquad\forall q_{h}\in Q_{h}({\omega_{\bm{a}}}).$ (24)
If $\mathcal{F}^{\star}=\emptyset$, i.e. none of the features in $\mathcal{F}$ is included into the geometry, then it is possible to prove that, given the regularity assumptions on $f$, the flux $\bm{\sigma}_{h}$ reconstructed from $u_{h}^{\star}=u_{h}^{0}$ is such that $||\nabla\cdot\bm{\sigma}_{h}-f||_{K}=0$, $\forall K\in\mathcal{A}_{h}=\mathcal{T}_{h}^{0}$. Equation (24) can indeed be written, for any $\bm{a}\in\mathscr{N}_{h}$, as
$(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h}+c_{\bm{a}})_{{\omega_{\bm{a}}}}=({\psi_{\bm{a}}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{0},q_{h}+c_{\bm{a}})_{{\omega_{\bm{a}}}}\qquad\forall q_{h}\in Q_{h}({\omega_{\bm{a}}}).$ (25)
Let us now observe that, according to (9), we have that
$(\nabla{\psi_{\bm{a}}}\cdot\nabla u_{h}^{0},1)_{{\omega_{\bm{a}}}}=({\psi_{\bm{a}}}f,1)_{{\omega_{\bm{a}}}}+\langle{\psi_{\bm{a}}}g_{\mathrm{N}},1\rangle_{\partial\omega_{\bm{a}}^{\psi}\cap\Gamma_{\mathrm{N}}^{0}}$ (26)
while, by the divergence theorem,
$(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},1)_{{\omega_{\bm{a}}}}=\langle\bm{\sigma}_{h}^{\bm{a}}\cdot\bm{n}_{\omega_{\bm{a}}},1\rangle_{\partial\omega_{\bm{a}}^{\psi}\cap\Gamma_{\mathrm{N}}^{0}}.$ (27)
Since the Neumann condition is strongly imposed on $\partial\omega_{\bm{a}}^{\psi}\cap\Gamma_{\mathrm{N}}^{0}$, we can rewrite (25) as
$(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h})_{{\omega_{\bm{a}}}}=({\psi_{\bm{a}}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{0},q_{h})_{{\omega_{\bm{a}}}},\quad\forall q_{h}\in Q_{h}({\omega_{\bm{a}}})$ (28)
Let us consider $K\in\mathcal{T}_{h}^{0}$ and let us denote its nodes by $\mathscr{N}_{h}(K)$. Since the polynomials in $Q_{h}({\omega_{\bm{a}}})$ are discontinuous, then (28) holds also for $q_{h}\in\mathcal{P}_{1}(K)$. We hence have that
$(\nabla\cdot\bm{\sigma}_{h},q_{h})_{K}$ $=\sum_{\bm{a}\in\mathscr{N}_{h}(K)}(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h})_{K}=\sum_{\bm{a}\in\mathscr{N}_{h}(K)}({\psi_{\bm{a}}}f,q_{h})_{K}=(f,q_{h})_{K}\quad\forall q_{h}\in\mathcal{P}_{1}(K),$ (29)
where we have also exploited the fact that ${\psi_{\bm{a}}}$ is a partition of unity and hence $\sum_{\bm{a}\in\mathscr{N}_{h}(K)}\nabla{\psi_{\bm{a}}}=0$.
In the second case, when $\mathcal{F}^{\star}\neq\emptyset$, the flux reconstructed from $u_{h}^{\star}$ is instead a weakly equilibrated flux reconstruction, due to the error introduced by the weak imposition of the Neumann boundary condition on ${\gamma^{\star}}$. Indeed $||\bm{\sigma}_{h}\cdot\bm{n}+g||_{{\gamma^{\star}}}$ will not be zero, and also the mass balance may be polluted, ending up in some elements for which $||\nabla\cdot\bm{\sigma}_{h}-f||_{K\cap{\Omega^{\star}}}\neq 0$.
Let us start from the case $p=0$, i.e. from the asymmetric version of the saddle-point problem (23). For any $q_{h}\in Q_{h}({\omega_{\bm{a}}})$, equation (24) can be rewritten as
@@
which holds also for $q_{h}\in\mathcal{P}_{1}(K)$, since the polynomials in $Q_{h}({\omega_{\bm{a}}})$ are discontinuous. We hence have that for any $K\in\mathcal{A}_{h}$ and for any $q_{h}\in\mathcal{P}_{1}(K)$
$(\nabla\cdot\bm{\sigma}_{h},q_{h})_{K\cap{\Omega^{\star}}}=\sum_{\bm{a}\in\mathscr{N}_{h}(K)}(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h})_{K\cap{\Omega^{\star}}}$ $=(f,q_{h})_{K\cap{\Omega^{\star}}}-\sum_{\bm{a}\in\mathscr{N}_{h}(K)}(\bm{\sigma}_{h}^{\bm{a}}\cdot\bm{n}+{\psi_{\bm{a}}}g,c_{\bm{a}})_{\gamma_{\bm{a}}^{\star}}$ (34)
Things are instead different if $p=1$. Indeed, for any $q_{h}\in Q_{h}({\omega_{\bm{a}}})$ equation (24) can be rewritten as
$(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h}+c_{\bm{a}})_{{\omega}_{\bm{a}}^{\star}}-\langle\bm{\sigma}_{h}^{\bm{a}}\cdot\bm{n}+{\psi_{\bm{a}}}g,q_{h}+c_{\bm{a}}\rangle_{\gamma_{\bm{a}}^{\star}}=({\psi_{\bm{a}}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{\star},q_{h}+c_{\bm{a}})_{{\omega}_{\bm{a}}^{\star}}$ (35)
while (31) and (32) are still holding. This means that (35) simply reduces to
$(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h})_{{\omega}_{\bm{a}}^{\star}}-\langle\bm{\sigma}_{h}^{\bm{a}}\cdot\bm{n}+{\psi_{\bm{a}}}g,q_{h}\rangle_{\gamma_{\bm{a}}^{\star}}=({\psi_{\bm{a}}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{\star},q_{h})_{{\omega}_{\bm{a}}^{\star}}\quad\forall q_{h}\in Q_{h}({\omega_{\bm{a}}}),$
and that, for any $K\in\mathcal{A}_{h}$ and $q_{h}\in\mathcal{P}_{1}(K)$
$(\nabla\cdot\bm{\sigma}_{h},q_{h})_{K\cap{\Omega^{\star}}}$ $=\sum_{\bm{a}\in\mathscr{N}_{h}(K)}(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h})_{K\cap{\Omega^{\star}}}=(f,q_{h})_{K\cap{\Omega^{\star}}}+\langle\bm{\sigma}_{h}\cdot\bm{n}+g,q_{h}\rangle_{\gamma_{K}^{\star}}.$ (36)
Let us remark that, in (34) the mass is not completely balanced in all the elements belonging to a cut patch, i.e. a patch ${\omega_{\bm{a}}}$ such that $|\gamma_{\bm{a}}^{\star}|\neq 0$. On the contrary, in (36), the mass is not balanced only in the elements for which $|\gamma_{K}^{\star}|\neq 0$, i.e. only for $K\in\mathcal{G}_{h}$. This means that, for $p=1$, and since we have assumed that $f\in Q_{h}(\Omega_{0})$,
$||\nabla\cdot\bm{\sigma}_{h}-f||_{K}=0\quad\forall K\in\mathcal{A}_{h}\setminus\mathcal{G}_{h}.$ (37)
@@
It is well known that the weak imposition of essential boundary conditions in CutFEM may require stabilization in presence of very small cuts [BURMAN_GHOST]. Let $v:\Omega_{{}_{\mathcal{T}}}\rightarrow\mathbb{R}$ be smooth enough and let us define the jump as
@@
$\mathscr{E}_{h,\bm{a}}:=\{e\in\mathscr{E}_{h}\text{ such that }~{}\exists K\in\mathcal{A}_{h},~{}K\subset{\omega_{\bm{a}}}:~{}e\subset\partial K\},\vspace{0.3cm}$
@@
and we define the bilinear forms ([BURMAN_GHOST, BurmanGeometry]) $\bm{j}_{h}:\bm{M}_{h}^{\bm{a}}\times\bm{M}_{h}^{\bm{a}}\rightarrow\mathbb{R}$, $j_{h}:Q_{h}^{\bm{a}}\times Q_{h}^{\bm{a}}\rightarrow\mathbb{R}$ such that
$\bm{j}_{\bm{a}}(\bm{\rho}_{h},\bm{v}_{h}):=\beta_{1}\sum_{e\in\mathscr{E}_{h,\bm{a}}^{\mathcal{G}}}\Big{[}h_{\bm{a}}(\llbracket\bm{\rho}_{h}\rrbracket_{e},\llbracket\bm{v}_{h}\rrbracket_{e})_{e}+h_{\bm{a}}^{3}(\llbracket\nabla\bm{\rho}_{h}\cdot\bm{n}_{e}\rrbracket_{e},\llbracket\nabla\bm{v}_{h}\cdot\bm{n}_{e}\rrbracket_{e})_{e}\Big{]}$ (38)
$j_{\bm{a}}(\mu_{h},q_{h}):=\beta_{2}\sum_{e\in\mathscr{E}_{h,\bm{a}}^{\mathcal{G}}}\Big{[}\frac{1}{h_{\bm{a}}}(\llbracket\mu_{h}\rrbracket_{e},\llbracket q_{h}\rrbracket_{e})_{e}+h_{\bm{a}}(\llbracket\nabla\mu_{h}\cdot\bm{n}_{e}\rrbracket_{e},\llbracket\nabla q_{h}\cdot\bm{n}_{e}\rrbracket_{e})_{e}\Big{]},$ (39)
with $\beta_{1},\beta_{2}>0$ and $\bm{n}_{e}$ denoting the unit normal vector to $e$, whose orientation is fixed. The stabilized version of Problem (23) then reads as: find $(\bm{\sigma}_{h}^{\bm{a}},\lambda_{h}^{\bm{a}})\in\bm{M}_{h}^{\bm{a}}\times Q_{h}^{\bm{a}}$ such that
$\begin{cases}m_{\bm{a}}(\bm{\sigma}_{h}^{\bm{a}},\bm{v}_{h})-b_{\bm{a}}^{1}(\bm{v}_{h},\lambda_{h}^{\bm{a}})+\bm{j}_{\bm{a}}(\bm{\sigma}_{h},\bm{v}_{h})=L_{\bm{a}}(\bm{v}_{h})&\forall\bm{v}_{h}\in\bm{M}_{h}^{\bm{a}}\\[6.00006pt] \hskip 62.59596ptb_{\bm{a}}^{p}(\bm{\bm{\sigma}_{h}^{\bm{a}}},q_{h})-j_{\bm{a}}(\lambda_{h}^{\bm{a}},q_{h})=R_{\bm{a}}^{p}(q_{h})&\forall q_{h}\in Q_{h}^{\bm{a}}.\end{cases}$ (40)
The scalar stabilization term introduced in the second equation is however likely to further pollute mass balance, not only on cut elements but on all the cut patches. Using a projection-based stabilization only on badly cut patches (see for example [puppi2021cut]) would allow to reduce the number of elements on which this additional error is introduced, but it might also affect the monotonicity of the estimator proposed in the next section. Indeed, assuming to use the estimator in an adaptive procedure, badly cut elements are very likely to appear and disappear as the mesh changes, introducing extra contributions to the estimator only at certain steps. Divergence preserving techniques have been proposed in [Zahedi_divFREE], requiring however the evaluation of higher order derivatives, and hence increasing the computational effort required to compute $\bm{\sigma}_{h}$.
Let us recall that $\bm{\sigma}_{h}$ is not the primal variable of our problem: it is computed only with the aim of building the estimator, so it is of interest to keep its computation as cheap as possible. For this reason, badly cut elements could also be simply neglected, allowing to avoid the case in which an element that has barely an active part, strongly affects the total estimator only due to stabilization. Further remarks on this aspect can be found in Section 6.
@@
In the present work we decided to focus on the negative feature case and on the challenges posed by the presence of cut elements. In the case of a positive feature included into the partially defeatured geometry, the equilibrated flux reconstruction in the feature and in the rest of the domain could be computed separately, provided that proper coupling conditions are defined at the interface.
What is interesting to remark is that, as reported in [BCGVV] (see also [BCV2022]), positive features with complex shapes are often treated by introducing a bounding box having the simplest possible shape, meshing the bounding box instead of the actual feature and computing the flux reconstruction on that mesh, which is blind to the feature boundary. Since a positive feature is actually a negative feature from its bounding box perspective, all the analysis that was carried out until now can easily be exploited.
@@ 4 A posteriori error analysis
In this section we propose a reliable estimator for the error $||\nabla(u-u_{h}^{\star})||_{\Omega}$, based on the (weakly) equilibrated flux defined above.
For $F\in\tilde{\mathcal{F}}$ let ${\tilde{\gamma}_{F}}:=\tilde{\gamma}\cap\partial F$ and $\tilde{\gamma}_{{0,F}}=\tilde{\gamma}_{0}\cap F$, so that $\tilde{\gamma}=\bigcup_{F\in\tilde{\mathcal{F}}}{\tilde{\gamma}_{F}}$ and $\tilde{\gamma}_{0}=\bigcup_{F\in\tilde{\mathcal{F}}}{\tilde{\gamma}_{0,F}}$. For every $F\in\tilde{\mathcal{F}}$ let ${\bm{\sigma}}|_{{F}}\in\bm{H}(\text{div},F)$ be the solution of
$\begin{cases}\nabla\cdot{\bm{\sigma}}|_{{F}}=f&\text{ in }F\\[3.00003pt] {\bm{\sigma}}|_{{F}}\cdot\bm{n_{0}}=-g_{0}&\text{ on }\tilde{\gamma}_{0,F}\\[3.00003pt] {\bm{\sigma}}|_{{F}}\cdot\bm{n_{F}}=-\bm{\sigma}\cdot\bm{n}&\text{ on }\tilde{\gamma}_{F},\end{cases}$ (41)
where $\bm{\sigma}\in\bm{H}(\text{div},{\Omega^{\star}})$ is a generic flux satisfying (17). Vectors $\bm{n_{F}}=-\bm{n}$ and $\bm{n_{0}}$ are the unitary outward normals of $F$ on $\tilde{\gamma}_{F}$ and $\tilde{\gamma}_{0,F}$, respectively. Omitting, by an abuse of notation, the explicit restriction of $\bm{\sigma}$ to $F$, we define for every $F\in\tilde{\mathcal{F}}$, the quantity
@@
$\mathcal{E}_{F}^{2}:=|{\tilde{\gamma}_{F}}|\big{|}\big{|}d_{F}^{h}-\overline{d_{F}^{h}}^{\tilde{\gamma}_{F}}\big{|}\big{|}_{{\tilde{\gamma}_{F}}}^{2}+c_{{\tilde{\gamma}_{F}}}^{2}|{\tilde{\gamma}_{F}}|^{2}\big{|}\overline{d_{F}}^{\tilde{\gamma}_{F}}\big{|}^{2},$ (42)
@@
$\mathcal{E}_{\mathrm{div}}^{K}:=h_{K}||f-\nabla\cdot\bm{\sigma}_{h}||_{K\cap{\Omega^{\star}}}\qquad\mathcal{E}_{\sigma}^{K}:=||\bm{\sigma}_{h}+\nabla u_{h}^{\star}||_{K\cap{\Omega^{\star}}}.$ (43)
$\mathcal{E}_{\mathrm{g}}^{K}:=h_{K}^{\frac{1}{2}}(||g+\bm{\sigma}_{h}\cdot\bm{n}||_{\gamma_{K}^{\star}}^{2}\big{)}^{\frac{1}{2}}$ (44)
We now state and prove the proposition establishing our a posteriori bound under the following technical assumption ([BCV2022_arxiv]):
@@
Let $u$ be the solution to (4) and $u_{h}^{\star}$ the solution to (16). Then, choosing $p=1$ in (23), and under Assumption 2, there exist three constants $\alpha_{1},\alpha_{2},\alpha_{3}>0$ independent on the mesh size $h$ and on the sizes of the features, such that
$||\nabla(u-u_{h}^{\star})||_{\Omega}\leq\Big{(}\sum_{K\in\mathcal{G}_{h}}\alpha_{1}(\mathcal{E}_{\mathrm{div}}^{K})^{2}\Big{)}^{\frac{1}{2}}+\Big{(}\sum_{K\in\mathcal{G}_{h}}\alpha_{2}(\mathcal{E}_{\mathrm{g}}^{K})^{2}\Big{)}^{\frac{1}{2}}+\Big{(}\sum_{K\in\mathcal{A}_{h}}(\mathcal{E}_{\sigma}^{K})^{2}\Big{)}^{\frac{1}{2}}+\Big{(}\sum_{F\in\tilde{\mathcal{F}}}\alpha_{3}\mathcal{E}_{F}^{2}\Big{)}^{\frac{1}{2}}$
@@
Let $v\in H_{0,\Gamma_{\mathrm{D}}}(\Omega)$. Adding and subtracting $(\bm{\sigma}_{h},\nabla v)_{\Omega}$, and exploiting (4) we have that
@@
Let $\mathbf{E}^{0}:H_{0,\Gamma_{\mathrm{D}}}(\Omega)\rightarrow H_{0,\Gamma_{\mathrm{D}}}(\Omega_{0})$ be a generalized Stein extension operator, such that $\forall v\in H_{0,\Gamma_{\mathrm{D}}}(\Omega)$
$||\nabla\mathbf{E}^{0}(v)||_{\Omega_{0}}\lesssim||\nabla v||_{\Omega}.$ (46)
For the construction of $\mathbf{E}^{0}$ we refer the reader to [Sauter], in which such extension operators are built for a large class of domains, and in particular for domains characterized by separated geometrical details of smaller scale. Let $v_{0}=\mathbf{E}^{0}(v)\in H_{0,\Gamma_{\mathrm{D}}}(\Omega_{0})$, and let us rewrite (45) as:
@@
We now introduce a Scott-Zhang type interpolation operator $I_{h}:H_{0,\Gamma_{\mathrm{D}}}^{1}(\Omega_{0})\rightarrow V_{h}^{0}(\Omega_{0})$ such that, $\forall v_{0}\in H_{0,\Gamma_{\mathrm{D}}}^{1}(\Omega_{0})$
$\sum_{K\in\mathcal{T}_{h}^{0}}h_{K}^{-2}||v_{0}-I_{h}v_{0}||_{K}^{2}\lesssim||\nabla v_{0}||_{\Omega_{0}}^{2},\qquad\sum_{K\in\mathcal{T}_{h}^{0}}||\nabla I_{h}v_{0}||_{K}^{2}\lesssim||\nabla v_{0}||_{\Omega_{0}}^{2},$ (48)

### v2

@@ 3 Error estimator via equilibrated flux reconstruction
If $\mathcal{F}^{\star}=\emptyset$, i.e. none of the features in $\mathcal{F}$ is included into the geometry, then it is possible to prove that, given the regularity assumptions on $f$, the flux $\bm{\sigma}_{h}$ reconstructed from $u_{h}^{\star}=u_{h}^{0}$ is such that $\nabla\cdot\bm{\sigma}_{h}-f||_{K}=0$, $\forall K\in\mathcal{A}_{h}=\mathcal{T}_{h}^{0}$ (see, e.g., [17, Lemma 3.5]).
When $\mathcal{F}^{\star}\neq\emptyset$, the flux reconstructed from $u_{h}^{\star}$ is instead a weakly equilibrated flux reconstruction, due to the error introduced by the weak imposition of the Neumann boundary condition on $\gamma^{\star}$}. Indeed $\bm{\sigma}_{h}\cdot\bm{n}+g||_{{\gamma^{\star}}}$ will not be zero, and also the mass balance may be polluted, ending up in some elements for which $\nabla\cdot\bm{\sigma}_{h}-f||_{K\cap{\Omega^{\star}}}\neq 0$. For any $\bm{a}\in\mathscr{N}_{h}$ we denote by $\overline{q_{h}}^{\omega_{\bm{a}}}$ the average of $q_{h}\in Q_{h}({\omega_{\bm{a}}})$ on $\omega_{\bm{a}}$}. Setting $c_{\bm{a}}=-\overline{q_{h}}^{\omega_{\bm{a}}}$, we have that $q_{h}+c_{\bm{a}}\in Q_{h}^{\bm{a}}$. The second equation in (25) can hence be rewritten as
$(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h}+c_{\bm{a}})_{{\omega}_{\bm{a}}^{\star}}-\langle\bm{\sigma}_{h}^{\bm{a}}\cdot\bm{n}+{\psi_{\bm{a}}}g,q_{h}+c_{\bm{a}}\rangle_{\gamma_{\bm{a}}^{\star}}=({\psi_{\bm{a}}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{\star},q_{h}+c_{\bm{a}})_{{\omega}_{\bm{a}}^{\star}}.$ (26)
@@
$(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h})_{{\omega}_{\bm{a}}^{\star}}-\langle\bm{\sigma}_{h}^{\bm{a}}\cdot\bm{n}+{\psi_{\bm{a}}}g,q_{h}\rangle_{\gamma_{\bm{a}}^{\star}}=({\psi_{\bm{a}}}f-\nabla\psi_{\bm{a}}\cdot\nabla u_{h}^{\star},q_{h})_{{\omega}_{\bm{a}}^{\star}}\quad\forall q_{h}\in Q_{h}({\omega_{\bm{a}}}).$ (29)
Let us consider $K\in\mathcal{A}_{h}$ and let us denote its nodes by $\mathscr{N}_{h}(K)$. As the polynomials in $Q_{h}({\omega_{\bm{a}}})$ are discontinuous, then (29) holds also for $q_{h}\in\mathcal{P}_{r}(K)$. We hence have that $\forall q_{h}\in\mathcal{P}_{r}(K)$
$(\nabla\cdot\bm{\sigma}_{h},q_{h})_{K\cap{\Omega^{\star}}}=\sum_{\bm{a}\in\mathscr{N}_{h}(K)}(\nabla\cdot\bm{\sigma}_{h}^{\bm{a}},q_{h})_{K\cap{\Omega^{\star}}}=(f,q_{h})_{K\cap{\Omega^{\star}}}+\langle\bm{\sigma}_{h}\cdot\bm{n}+g,q_{h}\rangle_{\gamma_{K}^{\star}},$ (30)
where we have exploited the fact that $\psi_{\bm{a}}$} is a partition of unity, and hence $\sum_{\bm{a}\in\mathscr{N}_{h}(K)}\nabla{\psi_{\bm{a}}}=0$. This means that, since we have assumed that $f\in Q_{h}(\Omega_{0})$,
$\nabla\cdot\bm{\sigma}_{h}-f||_{K}=0\quad\forall K\in\mathcal{A}_{h}\setminus\mathcal{G}_{h}.$ (31)
@@
The weakly equilibrated flux reconstruction could be built also by considering asymmetric local patch problems instead of (25), essentially by omitting the integrals over $\gamma_{\bm{a}}^{\star}$ in the mass balance equation [32]. However, as shown in Appendix A, this would imply a mass imbalance on a larger set of elements, namely not only in the cut elements, but in all the the elements belonging to a cut patch ($\omega_{\bm{a}}$} such that $|\gamma_{\bm{a}}^{\star}|\neq 0$). For this reason the symmetric formulation (25) is preferred.
Remark 3.
It is well known that the weak imposition of essential boundary conditions in CutFEM may require stabilization in presence of very small cuts [12] . Let $v:\Omega_{{}_{\mathcal{T}}}\rightarrow\mathbb{R}$ be smooth enough and let us define the jump as
@@
$\mathscr{E}_{h,\bm{a}}:=\{e\in\mathscr{E}_{h}\text{ such that }~\exists K\in\mathcal{A}_{h},~K\subset{\omega_{\bm{a}}}:~e\subset\partial K\},$
@@
and we define the bilinear forms ([12, 11]) $\bm{j}_{h}:\bm{M}_{h}^{\bm{a}}\times\bm{M}_{h}^{\bm{a}}\rightarrow\mathbb{R}$, $j_{h}:Q_{h}^{\bm{a}}\times Q_{h}^{\bm{a}}\rightarrow\mathbb{R}$ such that
$\bm{j}_{\bm{a}}(\bm{\rho}_{h},\bm{v}_{h}):=\beta_{1}\sum_{e\in\mathscr{E}_{h,\bm{a}}^{\mathcal{G}}}\Big[h_{\bm{a}}(\llbracket\bm{\rho}_{h}\rrbracket_{e},\llbracket\bm{v}_{h}\rrbracket_{e})_{e}+{\sum_{j=1}^{r}h_{\bm{a}}^{2j+1}(\llbracket D_{\bm{n}_{e}}^{j}\bm{\rho}_{h}\rrbracket_{e},\llbracket D_{\bm{n}_{e}}^{j}\bm{v}_{h}\rrbracket_{e})_{e}}\Big]$ (32)
$j_{\bm{a}}(\mu_{h},q_{h}):=\beta_{2}\sum_{e\in\mathscr{E}_{h,\bm{a}}^{\mathcal{G}}}\Big[\frac{1}{h_{\bm{a}}}(\llbracket\mu_{h}\rrbracket_{e},\llbracket q_{h}\rrbracket_{e})_{e}+{\sum_{j=1}^{r}h_{\bm{a}}^{2j-1}(\llbracket D_{\bm{n}_{e}}^{j}\mu_{h}\rrbracket_{e},\llbracket D_{\bm{n}_{e}}^{j}q_{h}\rrbracket_{e})_{e}}\Big],$ (33)
where $\beta_{1},\beta_{2}>0$ and $D_{\bm{n}_{e}}^{j}$ is the normal derivative of order $j$, with $\bm{n}_{e}$ denoting the unit normal vector to $e$, whose orientation is fixed. The stabilized version of Problem (25) then reads as: find $(\bm{\sigma}_{h}^{\bm{a}},\lambda_{h}^{\bm{a}})\in\bm{M}_{h}^{\bm{a}}\times Q_{h}^{\bm{a}}$ such that
$\begin{cases}m_{\bm{a}}(\bm{\sigma}_{h}^{\bm{a}},\bm{v}_{h})-{b_{\bm{a}}}(\bm{v}_{h},\lambda_{h}^{\bm{a}})+\bm{j}_{\bm{a}}(\bm{\sigma}_{h},\bm{v}_{h})=L_{\bm{a}}(\bm{v}_{h})&\forall\bm{v}_{h}\in\bm{M}_{h}^{\bm{a}}\\ \hskip 62.59596pt{b_{\bm{a}}}(\bm{\bm{\sigma}_{h}^{\bm{a}}},q_{h})-j_{\bm{a}}(\lambda_{h}^{\bm{a}},q_{h})={R_{\bm{a}}}(q_{h})&\forall q_{h}\in Q_{h}^{\bm{a}}.\end{cases}$ (34)
The scalar stabilization term introduced in the second equation is however likely to further pollute mass balance, not only on cut elements but on all the cut patches. Using a projection-based stabilization only on badly cut patches (see for example [32]) would allow to reduce the number of elements on which this additional error is introduced, but it might also affect the monotonicity of the estimator proposed in the next section. Indeed, assuming to use the estimator in an adaptive procedure, badly cut elements are very likely to appear and disappear as the mesh changes, introducing extra contributions to the estimator only at certain steps. Divergence preserving techniques have been proposed in [21], requiring however the evaluation of higher order derivatives, and hence increasing the computational effort required to compute $\bm{\sigma}_{h}$. Let us recall that $\bm{\sigma}_{h}$ is not the primal variable of our problem: it is computed only with the aim of building the estimator, so it is of interest to keep its computation as cheap as possible. For this reason, badly cut elements could also be simply neglected, allowing to avoid the case in which an element that has barely an active part, strongly affects the total estimator only due to stabilization. Further remarks on this aspect can be found in Section 6.
@@
In the present work we decided to focus on the negative Neumann feature case and on the challenges posed by the presence of cut elements when essential boundary conditions need to be weakly imposed in the equilibrated flux reconstruction. Referring to the terminology introduced in [8], also positive features could be considered, in the sense of small additions of material at the boundary of the domain. In that case the equilibrated flux reconstruction in the positive feature and in the rest of the domain could be computed separately, provided that proper coupling conditions are defined at the interface. What is interesting to remark is that, as reported in [7] (see also [8]), positive features with complex shapes are often treated by introducing a bounding box having the simplest possible shape, meshing the bounding box instead of the actual feature and computing the flux reconstruction on that mesh, which is blind to the feature boundary. Since a positive feature is actually a negative feature from its bounding box perspective, all the analysis that was carried out until now can easily be exploited. In contrast, the a posteriori error analysis for the Dirichlet feature case is more involved, as it requires to work on Sobolev spaces with negative exponent. We refer to [40] for the derivation of the defeaturing component of the estimator in that case.
@@ 4 A posteriori error analysis
In this section we propose a reliable estimator for the error $\nabla(u-u_{h}^{\star})||_{\Omega}$, based on the (weakly) equilibrated flux defined above.
For $F\in\tilde{\mathcal{F}}$ let ${\tilde{\gamma}_{F}}:=\tilde{\gamma}\cap\partial F$ and $\tilde{\gamma}_{{0,F}}=\tilde{\gamma}_{0}\cap F$, so that $\tilde{\gamma}=\bigcup_{F\in\tilde{\mathcal{F}}}{\tilde{\gamma}_{F}}$ and $\tilde{\gamma}_{0}=\bigcup_{F\in\tilde{\mathcal{F}}}{\tilde{\gamma}_{0,F}}$. Let $\bm{\sigma}\in\bm{H}(\text{div},{\Omega^{\star}})$ be a generic flux satisfying (17). Then, for every $F\in\tilde{\mathcal{F}}$ we denote by ${\bm{\sigma}}|_{{F}}\in\bm{H}(\text{div},F)$ its restriction to $F$, satisfying
$\begin{cases}\nabla\cdot{\bm{\sigma}}|_{{F}}=f&\text{ in }F\\ {\bm{\sigma}}|_{{F}}\cdot\bm{n_{0}}=-g_{0}&\text{ on }\tilde{\gamma}_{0,F}\\ {\bm{\sigma}}|_{{F}}\cdot\bm{n_{F}}=-\bm{\sigma}\cdot\bm{n}&\text{ on }\tilde{\gamma}_{F}.\end{cases}$ (35)
Vectors $\bm{n_{F}}=-\bm{n}$ and $\bm{n_{0}}$ are the unitary outward normals of $F$ on $\tilde{\gamma}_{F}$ and $\tilde{\gamma}_{0,F}$, respectively. Omitting, by an abuse of notation, the explicit restriction of $\bm{\sigma}$ to $F$, we define for every $F\in\tilde{\mathcal{F}}$, the quantity
@@
[…]
@@
[…]
We now state and prove the proposition establishing our a posteriori bound under the following technical assumption ([10]):
@@
Let $u$ be the solution to (4), $u_{h}^{\star}$ the solution to (16), and let $\bm{\sigma}_{h}$ be a weakly equilibrated flux reconstructed from $\nabla u_{h}^{\star}$ by solving the patch local problems in (25). Then, under Assumption 2, there exist three constants $\alpha_{1},\alpha_{2},\alpha_{3}>0$ such that
[…]
All the constants $\alpha_{i}$, $i=1,2,3$ may depend on the spatial dimension $d=2,3$, on the Lipschitz character of $\Omega$ and on the polynomial degree $r$, with $\alpha_{1}$ and $\alpha_{2}$ depending also on the shape regularity of the mesh ${\mathcal{T}_{h}}^{0}$. They are all independent on the feature size, on the mesh size, and on how the feature boundaries cut the mesh.
@@
Let $v\in H^{{1}}_{0,\Gamma_{\mathrm{D}}}(\Omega)$. Adding and subtracting $(\bm{\sigma}_{h},\nabla v)_{\Omega}$, and exploiting (4) we have that
@@
Let $\mathbf{E}^{0}:H_{0,\Gamma_{\mathrm{D}}}^{{1}}(\Omega)\rightarrow H_{0,\Gamma_{\mathrm{D}}}^{{1}}(\Omega_{0})$ be a generalized Stein extension operator, such that
[…]
In [33] an operator $\mathbf{E}^{0}$ satisfying (40) is built for a large class of domains, and in particular for domains characterized by separated geometrical details of smaller scale. In this case the constant $C_{\mathrm{stein}}$ is proven to be independent from the number and from the size of the features in $\mathcal{F}$. Let $v_{0}=\mathbf{E}^{0}(v)\in H_{0,\Gamma_{\mathrm{D}}}^{{1}}(\Omega_{0})$, and let us rewrite (39) as:
@@
We now introduce a projection operator $I_{h}:H_{0,\Gamma_{\mathrm{D}}}^{1}(\Omega_{0})\rightarrow{Q_{h}(\Omega_{0})}$ such that, $\forall v_{0}\in H_{0,\Gamma_{\mathrm{D}}}^{1}(\Omega_{0})$
[…]